For the following neutral atoms, give the number of protons, neutrons, and electrons:

204 80Hg

129 53I

7935Br

137 55Cs

192 78Pt
Answer #1

Number of electrons=atomic number in a neutral atom

Number of protons=atomic number

Number of neutrons=mass number-atomic number

204Hg 80

Electrons=80

Protons=80

Neutrons=204-80=124

129I53

Protons=53

Electrons=53

Neutrons=129-53=76

79Br 35

Electrons=35

Protons=35

Neutrons=79-35=44

137Cs55

Electrons=55

Protons=55

Neutrons=137-55=82

192pt78

Electrons=78

Protons=78

Neutrons=192-78=114